# Leveraging a Cognitive Model to Measure Subjective Similarity of Human and GPT-4 Written Content

### v1

@@ 1 Introduction
When humans categorize textual information, such as when giving recommendations or learning to categorize documents, we often use our personal subjective concepts to complete the task. One example of this is giving a recommendation of a funny book to a fiend, which requires not only our own subjective conceptualization of humor, but also an understanding of the similarities and differences between ourselves and our friends. While humans perform this task with relative ease, recommendation systems Ko et al. (2022) and educational tools Nafea et al. (2019) typically do not have personalized measurements of subjective concepts Gazdar and Hidri (2020), potentially hindering their efficacy Pal et al. (2024).
@@
When these systems incorporate data from human judgements to determine subjective similarity, they typically do so by pooling together as many judgements from different people as they can, and aggregate their measurement Xia et al. (2015). This approach relies on machine learning based methods Shojaei and Saneifar (2021), which can be effective from a machine learning perspective, since more data can mean improved document similarity metrics on average over large datasets Kusner et al. (2015). Focusing on individuals annotations of documents has been explored in the context of domain specific knowledge such as biomedical research papers Brown and Zhou (2019), or for specific context like document summarizing Zhang et al. (2003). However to date little attention has been given to the notion of individualized metrics of similarity that account for biases and constraints specifically, which are highly relevant for educational contexts Chew and Cerbin (2021).
@@
In this work, we propose a method for providing personalized metrics of subjective concepts that can determine the similarity between sets of text, with additional applications in textual categorization and educational feedback. This is done by leveraging a cognitive model of human learning and decision making that can act as a digital twin to individuals, and predict their behavior and opinions on a wider set of stimuli. We focus specifically on students categorizing emails as being safe (ham) or dangerous (phishing) in a training setting to help users identify and defend against phishing email attacks. Our proposed method for providing personalized similarity metrics of documents is compared to alternative methods using a dataset of a phishing education task experiment that we additionally present in this work.
Other than this method, the main contribution of this work is in presenting a dataset that has three components. The first component of this dataset is human annotations of email documents that are either written by cybersecurity experts or a GPT-4 model. The second component of this dataset is the emails that are shown to these participants, which allows for interesting comparisons between emails written by GPT-4 and cybersecurity experts, as wells as human subject’s perceptions of these emails. The final part of this dataset is a set of conversations between human participants and a GPT-4o model providing feedback to students.
@@
In total this dataset represents 39230 human judgements from 433 participants making decisions while observing a set from 1440 GPT-4 or human generated emails, as well as 20487 messages between human participants and the GPT-4o teacher model. This full database as well as all experimentation code discussed later is made available on OSF [^1] https://osf.io/wbg3r/.
2 Phishing Email Categorization Dataset
The first component of this dataset is human behavioral experiment data from a study on human categorization of emails. This experiment compares human document annotation when categorizing emails as phishing (dangerous) or ham (safe). The conditions of this experiment varied depending on the email author (Human or GPT-4) and style (plain-text or GPT-4 stylized). We also performed a comparison of document annotation training, which varied whether the emails shown to participants were randomly selected, or chosen using an IBL model (IBL or Random). Finally, we compared the type of feedback given to participants between positive and negative point feedback and a natural language conversation with an GPT-4o chat-bot (Points or Written).
The second component of this dataset is the emails shown to participants, which were either written by human cybersecurity experts, a GPT-4 model working alone, or a combination of human and GPT-4 model work. 360 base emails written by human experts were used to form three additional versions of these base emails. These alternative versions included a ‘human-written gpt4-styled’ version that used the email body written by human experts, the ‘gpt4-written and gpt4-styled’ version that was fully rewritten by GPT-4, and the ‘gpt4-written plaintext-styled’ version that stripped the HTML and CSS styling applied by the GPT-4 model. These emails as well as the original prompts to generate them are included in the presented dataset.
The final component of this dataset is a set of conversations between human participants and a GPT-4o model prompted to teach the participant to identify phishing emails. In this experiment three out of the eight experimental conditions involved human participants discussing the emails that they were categorizing with a GPT-4o model. This model was prompted to serve as an educational tool and varied in the type of information that was included in these prompts across experimental conditions. These teacher-student conversations consist of 20487 messages sent between human participants and the GPT-4o model.
Participants were allowed to send up to two messages to the teacher, and had a requirement to send relevant messages. If a participant categorized an email incorrectly, they were required to send at least one message to the teacher model. However, the analysis of human behavior and the embeddings formed by emails in this work focuses on human responses to categorization, and not the conversations between human participants and the training GPT-4o model.
@@ 3 Background: Cognitive Model
The cognitive model used in this work to predict the subjective similarity of human participants decisions on unseen emails relies on Instance Based Learning Theory (IBLT) Gonzalez et al. (2003). One of the benefits of employing IBL models over alternatives like Reinforcement Learning is that they base their predictions on the full history of participant experience as well as the impact that limitations like memory size and decay can have on decision making.
@@ 4.1 Human Subjective Similarity
To compare methods for evaluating the similarity of email documents, we use as a target our own devised metric of human subjective similarity. To determine this value, we use the category of human participant annotations, their annotation confidence, and the speed of their annotation. For accuracy and confidence, a higher value in our human subjective similarity metric signifies that participants were more likely to categorize an emails as being a member of that group. For reaction time, a lower value indicates that the document is more immediately obviously a member of a group and thus has a higher similarity to other members of that group.
The reaction time and confidence weighted subjective similarity of an email $x$ is given by multiplying the probability of a human participant categorizing that email as category $c$ giving $cs(x|c)=p(c|x)r(c|x)c(c|x)$. where $p(c|x)$ is the probability of categorization, $r(c|x)$ is the reaction time normalized to between 0 and 1, and $c(c|x)$ is the confidence additionally normalized to between 0 and 1. The soft-max of this $cs(x|c)$ value is the resulting similarity metric, with the equation shown in the supplementary materials.
Figure 1: Semantic and human participant similarity for phishing (blue) and ham (orange) emails. Shaded region is a logistic regression classifier. The Kernel Density Estimate log probability score between each distribution is shown on the bottom right, higher is better.
@@ 4.2 Semantic Similarity
One method of measuring the similarity of documents is to employ semantic information contained in documents and compare the similarities and differences between documents in terms of these semantic categories. This has been done in the past in applications such as topic modeling Řehůřek and Sojka (2010), document annotation Pech et al. (2017), and calculating document similarity Qurashi et al. (2020).
In this dataset, semantic similarity can be calculated using the categorizations of email features that were originally made by the cybersecurity experts who created the base email dataset. These features are Link Mismatch, Offer, Urgent, Subject Suspicious, Request Credentials, and Sender Mismatch. These six categories allow for a semantic similarity metric based on the presence of task-relevant features that are used by participants to annotate these documents. Figure 1 plots these semantic similarity measures for each of the 1440 emails in our dataset, and compares the distribution of these similarities to our human subjective similarity metric.
@@
Figure 2: Cosine and human participant similarity for phishing (blue) and ham (orange) emails. Shaded region is a logistic regression classifier. The Kernel Density Estimate log probability score between each distribution is shown on the bottom right, higher is better.
@@ 4.3 Cosine Similarity
Cosine similarity is the most commonly used metric of similarity of word and document embeddings, with many applications from classification Park et al. (2020), recommendation systems Khatter et al. (2021), educational tutorial systems Wu et al. (2023), question answering Aithal et al. (2021), and more Patil et al. (2023). However, there are limitations to using cosine similarity such as in documents with high-frequency words Zhou et al. (2022), and the presence of false information Borges et al. (2019), both of which are concerns for phishing email education.
The cosine similarity metric is calculated using an embedding of size 3072 formed by the ‘text-embedding-3-large’ model, accessed through the OpenAI API and calculated for each document in the database, these document embeddings are additionally included in our presented dataset. The equation for cosine similarity is included in the supplementary materials. The cosine similarity of each email embedding is compared to the mean embedding of that category and shown in Figure 2, and compared to our metric of human subjective similarity. From this, we can see that on average the embeddings are calculated as being significantly more similar to each other compared to the subjective similarities of human participants. This results in a lower Kernel Density Estimate log probability score (-2097.40) between the two distributions compared to the semantic similarity metric.
Figure 3: Cosine and human participant similarity for phishing (blue) and ham (orange) emails. Shaded region is a logistic regression classifier. The Kernel Density Estimate log probability score between each distribution is shown on the bottom right, higher is better.
@@ 4.4 Weighted Cosine Similarity
Distance weighted cosine similarity is a common method employed in utilizing embeddings Li and Han (2013), which has been applied onto measuring similarity of online instruction in educational settings Lahitani et al. (2016), as well as several cybersecurity specific applications like ransomware detection Moussaileb et al. (2021), and inside attacker detection Khan et al. (2019). In this work, we employ weighted cosine similarities of embeddings formed from emails categorized as being either ham or phishing, and compare it to human subjective similarity judgements. This weighting is done by learning a weight transformation of size 3072, the same as the embedding size, which is applied onto the embedding prior to calculating the similarity. The results of this weighting are shown in Figure 3, which compares the average human participant subjective similarity and the weighted cosine similarity of email embeddings.
@@
Figure 4: Pruned cosine and human participant similarity for phishing (blue) and ham (orange) emails. Shaded region is a logistic regression classifier. The Kernel Density Estimate log probability score between each distribution is shown on the bottom right, higher is better.
@@ 4.5 Pruning Document Embeddings
Another method of comparison documents is embedding pruning, where embeddings are reduced in size based on feedback from human categorizations to better account for their subjective similarity Manrique et al. (2023). These approaches function by reducing the number of embedding values that are used in comparison, and are similar to the weighting method except with 0 or 1 values. We structured our embedding pruning method to select only the top 500 embedding values, representing just under 20% of the size of the embedding, as was done in Manrique et al. (2023). These top predictive embedding values are retained, while all other values are masked to 0. After this, cosine similarity can be calculated with the standard approach, resulting in the similarity shown in Figure 5. Compared to the weighted cosine similarity method, the pruned cosine similarity has roughly the same KDE log probability score.
Figure 5: Ensemble and human participant similarity for phishing (blue) and ham (orange) emails. Shaded region is a logistic regression classifier. The Kernel Density Estimate log probability score between each distribution is shown on the bottom right, higher is better.
@@ 5 Ensemble Similarity
The final comparison method is based on using an ensemble of each of the previous similarity metrics, weighted to maximize the similarity to the average of the human subjective similarity metrics. This approach has been applied to document matching for patent documents Yu et al. (2024), which requires the similarity of document embeddings be calculated to determine a match. This ensemble approach has the highest KDE log probability score of any individual method by itself, at a value of -812.23. Looking at the KDE distributions above and to the right of the scatter plot in 5 demonstrates the high similarity of the ensemble similarity metric (light blue and light orange) and the human participant similarity metric (blue and orange). While this method is effective at resulting in a similarity metric that closely matches the average over all participants, it still does not fit as well to individual participants, as will be shown in our proposed model.
Figure 6: IBIS and human participant similarity for phishing (blue) and ham (orange) emails. Shaded region is a logistic regression classifier. The Kernel Density Estimate log probability score between each distribution is shown on the bottom right, higher is better.
Figure 7: Top performing similarity metrics and individual participant similarity for phishing and ham emails. Shaded region is a logistic regression classifier. The lower value is the individual KDE score, with averages of all participants shown in Table 1
@@ 6 Instance-Based Individualized Similarity (IBIS)
Predictions of Instance-Bases Individual Similarity are done using an IBL model that is currently serving as a digital twin with the same experience as an individual participant. Using this we determine the value that the IBL model assigns to predicting a category $c$ as $V_{k}(c|x)$, or the value the IBL model assigns to choosing option $c$ as the category of document $x$. Then, we can divide this value by the same categorization value assigned to each alternative categorization of the same document. This results in the IBIS metric which can be calculated after each decision is made by a participant, pseudo-code for the IBIS algorithm is included in the supplementary materials. The code-base for the IBIS method including all comparison methods, data, and scripts to generate similarity measures and figures is made available222github.com/TylerJamesMalloy/cognitive-similarity.

### v2

@@ 1 Introduction
When humans categorize textual information, such as when giving recommendations or learning to categorize documents, we often use our personal subjective concepts to complete the task. One example of this is giving a recommendation of a funny book to a fiend, which requires not only our own subjective conceptualization of humor, but also an understanding of the similarities and differences between ourselves and our friends. While humans perform this task with relative ease, recommendation systems (Ko et al., 2022) and educational tools (Nafea et al., 2019) typically do not have personalized measurements of subjective concepts (Gazdar and Hidri, 2020), potentially hindering their efficacy (Pal et al., 2024).
@@
When these systems incorporate data from human judgements to determine subjective similarity, they typically do so by pooling together as many judgements from different people as they can, and aggregate their measurement (Xia et al., 2015). This approach relies on machine learning based methods (Shojaei and Saneifar, 2021), which can be effective from a machine learning perspective, since more data can mean improved document similarity metrics on average over large datasets (Kusner et al., 2015). Focusing on individuals annotations of documents has been explored in the context of domain specific knowledge such as biomedical research papers (Brown and Zhou, 2019), or for specific context like document summarizing (Zhang et al., 2003).
However to date little attention has been given to the notion of individualized metrics of similarity that account for biases and constraints specifically, which are highly relevant for educational contexts (Chew and Cerbin, 2021). Related to the domain of this work in particular, recent work has demonstrated a broad range of human opinions and levels of trust associated with cybersecurity concepts such as Trusted Execution Environments (Carreira et al., 2024). This highlights the need for individualized metrics that take into account experience in training tasks such as the anti-phishing training dataset used in this work.
@@
In this work, we propose a method for providing personalized metrics of subjective concepts that can determine the similarity between sets of text, with additional applications in selecting educational examples and providing natural language feedback. This is done by leveraging a cognitive model of human learning and decision making that can act as a digital twin to individuals, and predict their behavior and opinions on a wider set of stimuli. We focus specifically on students categorizing emails as being safe (ham) or dangerous (phishing) in a training setting to help users identify and defend against phishing email attacks. Our proposed method for providing personalized similarity metrics of documents is compared to alternative methods using a dataset of a phishing education task experiment that we additionally present in this work.
@@
The dataset of human annotations of emails as being either ham or phishing is described in (Malloy et al., 2024) and was made publicly available on OSF [^1] https://osf.io/wbg3r/. This dataset consists of human annotations of email documents that are either written by cybersecurity experts or a GPT-4 model, the emails shown to participants, and conversations between human participants and a GPT-4o model providing feedback to students. In total this dataset represents 39230 human judgements from 433 participants making decisions while observing a set from 1440 GPT-4 or human generated emails, as well as 20487 messages between human participants and the GPT-4o teacher model.
This type of learning task represents a serious challenge for traditional methods of adjusting document embedding similarity metrics to conform to human behavior, such as embedding pruning (Manrique et al., 2023) or embedding weighting (Onan, 2021). This is because these approaches typically rely on a large amount of annotations collected from many participants who are expected to have the same knowledge level throughout the annotation process. Instead, we are interested in measuring the subjective similarity of documents as participants learn the document annotation task in a training setting. To do this, we employ a cognitive model that can predict the learning trajectory of each individual participant as they learn to correctly annotate these documents.
@@ 2 Background: Cognitive Model
The cognitive model used in this work to predict the subjective similarity of human participants decisions on unseen emails relies on Instance Based Learning Theory (IBLT) (Gonzalez et al., 2003). One of the benefits of employing IBL models over alternatives like Reinforcement Learning is that they base their predictions on the full history of participant experience as well as the impact that limitations like memory size and decay can have on decision making.
@@ 2.1 Activation
2.2 Similarity Measure
The definition of the similarity measure $S_{ij}$ is highly influential in the behavior of the IBL model, as it determines which instances from memory are drawn from to predict utility. In simple binary choice tasks without attributes (Gonzalez and Dutt, 2011; Lejarraga et al., 2012), the similarity metric can be defined as the equality function $S_{ij}=1$ if $i==j$ else 00. In more complex domains such as the phishing email identification task used in this work, one approach is to use the embeddings of emails to compare the similarity of instances, and rely on the cosine similarity metric to compute the similarity of instances in memory (Malloy and Gonzalez, 2024). The model presented in this work relies on an initial baseline similarity metric, the standard cosine similarity, to then build more individual specific metrics of similarity.
@@ 2.4 Blended Value
These blended values are used to determine the action $a_{t+1}$ selected by the model at the next time step.
$a_{t+1}=\max_{k\in K}V_{k}(t)$ (4)
In standard IBL models, this action can be used in simulations to allow the model to gain experience in a given task. In model tracing, which is used in the method proposed in this work, the memory of instances is made up of the past observations and decisions of the participant, with the action representing a prediction of their future behavior.
3 Phishing Email Categorization Dataset
The first component of this dataset is human behavioral experiment data from a study on human categorization of emails. This experiment compared human document annotation when categorizing emails as phishing (dangerous) or ham (safe). The conditions of this experiment varied depending on the email author (Human or GPT-4) and style (plain-text or GPT-4 stylized). There was also a comparison of the method of selecting emails to show to participants, either randomly selected, or chosen using an IBL model (IBL or Random). Finally, we compared the type of feedback given to participants between positive and negative point feedback and a natural language conversation with an GPT-4o chat-bot (Points or Written).
The second component of this dataset is the emails shown to participants, which were either written by human cybersecurity experts, a GPT-4 model working alone, or a combination of human and GPT-4 model work. 360 base emails written by human experts were used to form three additional versions of these base emails. These alternative versions included a ‘human-written gpt4-styled’ version that used the email body written by human experts, the ‘gpt4-written and gpt4-styled’ version that was fully rewritten by GPT-4, and the ‘gpt4-written plaintext-styled’ version that stripped the HTML and CSS styling applied by the GPT-4 model. These emails as well as the original prompts to generate them are included in dataset on OSF.
@@ 4.1 Human Subjective Similarity
LLM embeddings have been suggested as a method of measuring human similarity judgements (Bhatia and Aka, 2022), while also capturing the wide range of individuals similarity measures. Additionally, comparisons of LLM behavior have also demonstrated human-like variability (Bhatia, 2024), suggesting these embeddings could be useful for capturing the variety of human similarity judgements. Cognitive models that rely on representations of information from GAI models have been shown to adequately account for the wide range of human behavior (Mitsopoulos et al., 2023).
However, for these methods to function properly there must be a connection between the way that similarity is measured in humans and GAI models. Previous applications in applying visual GAI models onto representing decision-making tasks in humans relied on the close connection to these model representations and human representations (Higgins et al., 2016, 2021). For this reason, we devised a metric of human subjective similarity that takes into account the confidence of document categorization as well as the time it takes participants to categorize documents.
Figure 1: Human participant similarity measure for all 1440 phishing (blue) and ham (orange) emails. Shaded region is a logistic regression.
To determine the human subjective similarity measure, we use the category of human participant annotations, their annotation confidence, and the speed of their annotation. For accuracy and confidence, a higher value in our human subjective similarity metric signifies that participants were more likely to categorize an emails as being a member of that group, and more confident in their categorization. For reaction time, a lower value indicates that the document is more immediately obviously a member of a group and thus has a higher similarity to other members of that group. The result is a value that is difficult for a standard similarity metric to account for, as the annotations made in this dataset occurred in a learning setting where earlier trials had less accuracy, which also impacted reaction time and confidence.
The reaction time and confidence weighted subjective similarity of an email $x$ is given by multiplying the probability of a human participant categorizing that email as category $c$ giving $cs(x|c)=p(c|x)r(c|x)c(c|x)$. where $p(c|x)$ is the probability of categorization, $r(c|x)$ is the reaction time normalized to between 0 and 1, and $c(c|x)$ is the confidence additionally normalized to between 0 and 1. The soft-max of this $cs(x|c)$ value is the resulting similarity metric, with the equation shown in the supplementary materials [^2] https://osf.io/wbg3r/.
$HS(x,x^{\prime})=\dfrac{cs(x|c)cs(x^{\prime}|c)}{\sum_{c^{\prime}\in C}cs(x|c)\sum_{c^{\prime}\in C}cs(x^{\prime}|c^{\prime})}$ (5)
Figure 1 shows the average human similarity measures for each of the 1440 emails in the dataset. The human ham and human phishing similarities are calculated according to Equation 5 by averaging the accuracy, reaction time, and confidence across all participants in the dataset. It is also possible to calculate this subjective similarity for an individual only using the documents that subject categorized. We next compare similarity measures in their ability to capture individual human subjective similarity.
Figure 2: Semantic and human participant similarity for phishing (blue) and ham (orange) emails. Shaded region is a logistic regression. The Kernel Density Estimate log probability score between each distribution is shown on the bottom right, higher is better.
Figure 3: Cosine and human participant similarity for phishing (blue) and ham (orange) emails. Shaded region is a logistic regression. The Kernel Density Estimate log probability score between each distribution is shown on the bottom right, higher is better.
@@ 4.2 Semantic Similarity
One method of measuring the similarity of documents is to employ semantic information contained in documents and compare the similarities and differences between documents in terms of these semantic categories. This has been done in the past in applications such as topic modeling (Řehůřek and Sojka, 2010), document annotation (Pech et al., 2017), and calculating document similarity (Qurashi et al., 2020).
In this dataset, semantic similarity can be calculated using the categorizations of email features that were originally made by the cybersecurity experts who created the base email dataset. These features are Link Mismatch, Offer, Urgent, Subject Suspicious, Request Credentials, and Sender Mismatch. Figure 2 plots these semantic similarity measures for each of the 1440 emails in our dataset, and compares the distribution of these similarities to our human subjective similarity metric.
@@
Figure 4: Cosine and human participant similarity for phishing (blue) and ham (orange) emails. Shaded region is a logistic regression. The Kernel Density Estimate log probability score between each distribution is shown on the bottom right, higher is better.
@@ 4.3 Cosine Similarity
Cosine similarity is the most commonly used metric of similarity of word and document embeddings, with many applications from classification (Park et al., 2020), recommendation systems (Khatter et al., 2021), educational tutorial systems (Wu et al., 2023), question answering (Aithal et al., 2021), and more (Patil et al., 2023). However, there are limitations to using cosine similarity such as in documents with high-frequency words (Zhou et al., 2022), and the presence of false information (Borges et al., 2019), both of which are concerns for phishing email education.
The cosine similarity metric is calculated using an embedding of size 3072 formed by the ‘text-embedding-3-large’ model, accessed through the OpenAI API, these document embeddings are additionally included in our presented dataset. The cosine similarity of each email embedding is compared to the mean embedding of that category and shown in Figure 3, and compared to our metric of human subjective similarity. From this, we can see that on average the embeddings are calculated as being significantly more similar to each other compared to the subjective similarities of human participants. This results in a lower Kernel Density Estimate log probability score (-2097.40) between the two distributions compared to the semantic similarity metric.
@@ 4.4 Weighted Cosine Similarity
Distance weighted cosine similarity is a common method employed in utilizing embeddings (Li and Han, 2013), which has been applied onto measuring similarity of online instruction in educational settings (Lahitani et al., 2016), as well as several cybersecurity specific applications like ransomware detection (Moussaileb et al., 2021), and inside attacker detection (Khan et al., 2019). In this work, we employ weighted cosine similarities of embeddings formed from emails categorized as being either ham or phishing, and compare it to human subjective similarity judgements. This weighting is done by learning a weight transformation of size 3072, the same as the embedding size, which is applied onto the embedding prior to calculating the similarity. The results of this weighting are shown in Figure 4, which compares the average human participant subjective similarity and the weighted cosine similarity of email embeddings.
@@ 4.5 Pruning Document Embeddings
Figure 5: Pruned cosine and human participant similarity for phishing (blue) and ham (orange) emails. Shaded region is a logistic regression. The Kernel Density Estimate log probability score between each distribution is shown on the bottom right, higher is better.
Another method of comparison documents is embedding pruning, where embeddings are reduced in size based on feedback from human categorizations to better account for their subjective similarity (Manrique et al., 2023). These approaches function by reducing the number of embedding values that are used in comparison, and are similar to the weighting method except with 0 or 1 values. We structured our embedding pruning method to select only the top 500 embedding values, representing just under 20% of the size of the embedding, as was done in (Manrique et al., 2023). These top predictive embedding values are retained, while all other values are masked to 0. After this, cosine similarity can be calculated with the standard approach, resulting in the similarity shown in Figure 6. Compared to the weighted cosine similarity method, the pruned cosine similarity has roughly the same KDE log probability score.
@@ 5 Ensemble Similarity
Figure 6: Ensemble and human participant similarity for phishing (blue) and ham (orange) emails. Shaded region is a logistic regression. The Kernel Density Estimate log probability score between each distribution is shown on the bottom right, higher is better.
The final comparison method is based on using an ensemble of each of the previous similarity metrics, weighted to maximize the similarity to the average of the human subjective similarity metrics. This approach has been applied to document matching for patent documents (Yu et al., 2024), which requires the similarity of document embeddings be calculated to determine a match. This ensemble approach has the highest KDE log probability score of any individual method by itself, at a value of -812.23. Looking at the KDE distributions above and to the right of the scatter plot in 6 demonstrates the high similarity of the ensemble similarity metric (light blue and light orange) and the human participant similarity metric (blue and orange). While this method is effective at resulting in a similarity metric that closely matches the average over all participants, it still does not fit as well to individual participants, as will be shown in our proposed model.
@@ 6 Instance-Based Individualized Similarity (IBIS)
Figure 7: IBIS and human participant similarity for phishing (blue) and ham (orange) emails. Shaded region is a logistic regression. The Kernel Density Estimate log probability score between each distribution is shown on the bottom right, higher is better.
@@
Predictions of Instance-Bases Individual Similarity are done using an IBL model that is currently serving as a digital twin with the same experience as an individual participant. Using this we determine the value that the IBL model assigns to predicting a category $c$ as $V_{k}(c|x)$, or the value the IBL model assigns to choosing option $c$ as the category of document $x$. Then, we can divide this value by the same categorization value assigned to each alternative categorization of the same document. This results in the IBIS metric which can be calculated after each decision is made by a participant, pseudo-code for the IBIS algorithm, The code-base for the IBIS method including all comparison methods, data, and scripts to generate similarity measures and figures is made available333github.com/TylerJamesMalloy/cognitive-similarity.